Another similar problem to the one posted earlier, but with a different volume...

• I count 176 colonies on a plate onto which 100 ml of a 10^-5 dilution of my culture was plated.

•what was the concentration of my original culture?
Answer #1

So, this problem needs calculation of CFU/ml first.

Cfu/ml = Number of colonies/Vomume of sample plated*dilution factor

= 176/100*105 = 176/107 = 176*10-7 or 1.76*10-9 cfu/ml

To get the concentration of original sample, the cfu/ml has to be mutiplied with the dilution factor.

So, it becomes = 1.76*10-9 cfu/ml * 105 = 1.76*10-2 cfu/ml

Therefore, 1.76*10-2 cfu/ml is the concentration of original sample.

Hope it helps, Good luck !!!!
Please do press like :)